# Compact Kähler manifolds with quasi-positive holomorphic sectional curvature

### v3

@@ 2.3. Some properties of curvature tensor
Let $(M,\omega)$ be a compact Kähler manifold. Assume that $\mathcal{F}$ is an invertible sheaf of $\mathcal{O}(\Lambda^{p,0}M)$ which is pseudo-effective as a line bundle. Then there exist a finite covering $\{U_{\alpha}\}$ consists of coordinate neighborhoods and local holomorphic $(p,0)$-forms $\eta_{\alpha}$ generating $\mathcal{F}|_{U_{\alpha}}$, $\mathcal{F}$ is associated with the Čech cocycle $b_{\alpha\beta}$ in $\mathcal{O}_{M}^{*}(U_{\alpha})$ such that $\eta_{\beta}=b_{\alpha\beta}\eta_{\alpha}$ and there exists a singular Hermitian metric $h=e^{-\varphi}$ of $\mathcal{F}$ defined by a collection of plurisubharmonic functions $\varphi_{\alpha}\in\mathrm{PSH}(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$ such that $e^{-\varphi_{\beta}}=|b_{\alpha\beta}|^{2}e^{-\varphi_{\alpha}}$. We can define a globally bounded measurable function $\psi=e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}$ as in ([8]), which can be viewed as the hermitian metric ratio $\frac{|\eta_{\alpha}|_{g}}{|\eta_{\alpha}|_{h}}$. Since $\sqrt{-1}\partial\bar{\partial}$ of the plurisubharmonic function $\varphi_{\alpha}$ is a closed positive current,
$\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}})=e^{\varphi_{\alpha}}(\sqrt{-1}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha})$
is also a closed positive current on $U_{\alpha}$. This shows that $\partial\varphi_{\alpha}$ is $L^{2}$ with respect to the wight $e^{\varphi_{\alpha}}$ and $e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}$ has locally finite measure corfficients on $U_{\alpha}$. Hence
@@
This subsection is devoted to the proof of Proposition 3.1. Firstly, we will proof the following key identity for forms.
@@
$\omega^{n-p-1}$
$=$ $(\sqrt{-1})^{n-p-1}(n-p-1)!\sum\limits_{i_{1}<i_{2}<\cdots<i_{n-p-1}}dz^{i_{1}}\wedge d\bar{z}^{i_{1}}\wedge\cdots\wedge dz^{i_{n-p-1}}\wedge d\bar{z}^{i_{n-p-1}},$
@@
$(\sqrt{-1})^{p^{2}}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}$
$=$ $(\sqrt{-1})^{p+1}\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{i}\wedge d\bar{z}^{i}\wedge dz^{j_{1}}\wedge d\bar{z}^{j_{1}}\wedge\cdots\wedge dz^{j_{p}}\wedge d\bar{z}^{j_{p}}$
@@
where $\beta_{\alpha}=\Lambda^{p-1}(\frac{\eta_{\alpha}\wedge\overline{\eta_{\alpha}}}{(p-1)!})$. $e^{\varphi_{\alpha}}\beta_{\alpha}$ and $e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}$ are globally defined differential forms with bounded measurable cofficients. For clarity, the arguement will be divided into three steps.
@@
Although $\varphi_{\alpha}$ is in general singular, we can still do the integration by parts as in [12]. For reader’s convenience, we give a detailed explaination. Take a partitions of unity $\{\rho_{\alpha}\}$ subordinated to the finite cover $\{U_{\alpha}\}$. Since $\mathrm{supp}\rho_{\alpha}\subset\subset U_{\alpha}$, for any $\epsilon>0$ sufficiently small, there exists a sequence of smooth plurisubharmonic functions $\varphi_{\alpha,\epsilon}$ decreasing to $\varphi_{\alpha}$ on some open subset $V_{\alpha}$ such that $\mathrm{supp}\rho_{\alpha}\subset\subset V_{\alpha}\subset\subset U_{\alpha}$. Now the results of Bedford-Taylor (Lemma 2.4) applied to the uniformly bounded function $e^{\varphi_{\alpha,\epsilon}}$ on $V_{\alpha}$, imply that we have local weak covergence
@@
The last equality follows from the fact $\sum\limits_{\alpha}\rho_{\alpha}=1$. Combine with (3.9), we get
@@
Next, let us separately analysis each term on the right-hand side of integral inequality (3.11). Firstly, we consider the sign of first term of (3.11).
@@
For any $x_{0}\in U_{\alpha}$, we choose local holomorphic coordinates centered at $x_{0}$ such that
@@
at $x_{0}$. For simplicity, we omit the index $\alpha$ in the following calculation. We write $\eta=\sum\limits\eta_{I_{p}}dz^{I_{p}}$, recall (3.3) and (3.7) that we have
@@
By a direct calculation, we derive the following Bochner type formula in the sense of distribution,
@@
$\begin{split}e^{2\varphi}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}=&e^{2\varphi}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}}\\ \geq&e^{2\varphi}\frac{\kappa(x_{0})}{2}\left[(\sum\limits_{i}\beta_{i\bar{i}})^{2}+\sum\limits_{i}\beta_{i\bar{i}}^{2}\right]\\ =&\frac{\kappa(x_{0})}{2}e^{2\varphi}\left(|\eta|_{g}^{4}+|\beta|_{g}^{2}\right)\\ \geq&\frac{\kappa(x_{0})}{2}\psi^{2}\end{split}$ (3.16)
@@
¯$\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}})=\partial(e^{\varphi_{\alpha}}\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}}+(-1)^{p}\eta_{\alpha}\wedge\overline{\partial\eta_{\alpha}})$
@@
Hence $d\psi=0$ almost everywhere on $M$, which implies that $\psi=C$ for some constant $C\geq 0$ almost everywhere. We claim that $C>0$. Suppose not, we have $\psi=0$ almost everywhere on $U_{\alpha}$. Since $\varphi_{\alpha}\in Psh(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$, $\{\varphi_{\alpha}=-\infty\}$ has zero Lebesgue measure. This fact implies that $\eta_{\alpha}=0$ almost everywhere on $U_{\alpha}$. Then $\eta_{\alpha}\equiv 0$ on $U_{\alpha}$ due to the smoothness of $\eta_{\alpha}$, which contradicts that $\eta_{\alpha}$ generates $\mathcal{F}|_{U_{\alpha}}$. Therefore, $\varphi_{\alpha}<+\infty$ and $e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}=C>0$ implies that $|\eta_{\alpha}|^{2}_{g}>0$ and $\varphi_{\alpha}=\log C-\log|\eta_{\alpha}|^{2}_{g}$ is a smooth function on $U_{\alpha}$. The following calculation will be done in the smooth sense.
@@
Since $x_{0}\in M$ is arbitrary, we have ¯$\partial\bar{\partial}\varphi=0$ on $M$ and $D^{\prime}\eta+\eta\partial\varphi=0$ on $M$. ¯$\partial\bar{\partial}\varphi_{\alpha}=0$ implies that $(\mathcal{F},e^{-\varphi})$ is a flat line bundle. In each coordinate neighborhood $U_{\alpha}$, $\varphi_{\alpha}$ is pluriharmonic, then $\varphi_{\alpha}$ can be written $\varphi_{\alpha}=f_{\alpha}+\overline{f_{\alpha}}$ for some holomorphic function $f_{\alpha}\in\mathcal{O}(U_{\alpha})$ and
$D^{\prime}(e^{f_{\alpha}}\eta_{\alpha})=e^{f_{\alpha}}(D^{\prime}\eta_{\alpha}+\eta_{\alpha}\partial\varphi_{\alpha})=0$
where $e^{f_{\alpha}}\eta_{\alpha}$ is a local holomorphic section of $\mathcal{F}|_{U_{\alpha}}$. This shows that $\mathcal{F}$ is invariant by parallel transport induced by connection $D$ of $\Lambda^{p,0}M$. The proof is completed. ∎
4. Proof of Theorem 1.5 and Theorem 1.7
In this subsection, we proceed to show Theorem 1.5 and Theorem 1.7. The proof of Theorem 1.5 is based on Theorem 1.4 (1) and Campana-Demailly-Peternell’s critetion. The proof of Theorem 1.7 follows the arguement in [7, section 4] for $\operatorname{Ric}\geq 0$, demonstrating that factors of irreducible holonomy $H=U(n)$ are rationally connected manifolds. The difference is that factors of irreducible holonomy $H=SU(n)$ or $Sp(\frac{n}{2})$ can be excluded by statement (1) of Theorem 1.4.
@@
Since the holomorphic sectional curvature is nonnegative and there exists no nonzero truly flat tangent vector, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective by the statement (1) of Theorem 1.4. This fact implies that $H^{p,0}_{\bar{\partial}}(M)=0,\ \forall 1\leq p\leq n$, in particular, M is projective by Kodaira’s theorem ([21]). Combine with Campana-Demailly-Peternell’s criterion (Lemma 2.5), $M$ is rationally connected. ∎
@@
Let $(M,\omega)\simeq\prod(M_{l},\omega_{l})$ be the De Rham decomposition of $(M,\omega)$, induced by a decomposition of the holonomy representation in irreducible representations ([4]). Since the holonomy is contained in $U(n)$, all factors $(M_{l},\omega_{l})$ are Kähler manifolds with irreducible holonomy and holonomy group $H_{l}\subset U(n_{l})$, $n_{l}=\dim M_{l}$. The product structure implies that $\omega_{l}$ has nonnegative holomorphic sectional curvature. By Berger’s classification of holonomy groups ([4]), there are only three possibilities: $H_{l}=U(n_{l}),\ SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$, unless $M_{l}$ is a Hermitian symmetric space, then necessarily of compact type (such symmetric space were classified by Élie Cartan, they are rational homogeneous, hence rationally connected, their holonomy groups are also known, see. e.g. [5]).
Case 1: $H_{l}=SU(n_{l})$ or $Sp(\frac{n_{l}}{2})$. Then $M_{l}$ is a Calabi-Yau manifold or a holomorphic symplectic manifold (see e.g. [1]). Recall that a simply connected Calabi-Yau manifold $M_{l}$ is a compact simply connected projective manifold with $K_{M_{l}}\simeq O_{M_{l}}$ and $H^{0}(M_{l},\Omega_{M_{l}}^{p})=0$ for $0<p<\dim M_{l}$. A holomorphic symplectic manifold is a compact Kähler manifold $X$ which admits a holomorphic symplectic $(2,0)$-form (of maximal rank everywhere). In particular, $M_{l}$ admits a holomorphic $(n_{l},0)$-form that vanishes nowhere. Recall that we can define a nonnegative $(1,1)$-form $\beta=\Lambda^{p-1}({\frac{\eta\wedge\overline{\eta}}{n_{l}!}})$. For any point $x_{0}$ of $M_{l}$, we can choose some local holomorphic coordinates such that the coefficients of $w_{l}$ is identity and $\beta$ is diagonal at $x_{0}$. Locally, we can write $\eta=\frac{1}{n_{l}!}\sum\limits fdz^{1}\wedge\cdots dz^{n}$. It follows from the local representation (3.7) of $\beta$ that we have
$\beta=\frac{|\eta|^{2}}{n_{l}}\sum\limits_{i=1}^{n_{l}}dz^{i}\wedge d\overline{z}^{i}$
at $x_{0}$. The proof of Theorem 1.4 tells us that $\{\frac{\partial}{\partial z_{i}}\}_{i=1}^{n_{l}}$ are truly flat tangent vectors at $x_{0}$. Hence the Riemannian curvature tensor $\operatorname{Rm}$ of $\omega_{l}$ vanishes at $x_{0}$. Since $x_{0}$ is arbitrary, $\operatorname{Rm}$ vanishes everywhere, $M_{l}$ is a simply connected complete manifold with zero holomorphic sectional curvature. By classification of complex space forms, $M_{l}$ must be holomorphically isometric to complex Euclidean space $\mathbb{C}^{n_{l}}$, which contradicts the fact that $M_{l}$ is compact.
Case 2: $H_{l}=U(n_{l})$. Since the natural representation of $U(n_{l})$ on $(\mathbb{C}^{n_{l}})^{\otimes p}$ has no invariant line on which $U(n_{l})$ induces a trivial action, there cannot exist any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M_{i})$, $1\leq p\leq n$ such that $\mathcal{F}$ is invariant by parallel transport induced by the connection $D$. Combine with statement (2) of Theorem 1.4, any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective. Therefore $M_{l}$ is rationally connected by Campana-Demailly-Peternell’s criterion.
In summary, each $(M_{l},\omega_{l})$ is a rationally connected manifold with nonnegative holomorphic sectional curvature and irreducible holonomy (holonomy $U(n_{l})$ unless $M_{l}$ is a hermitian symmetric space of compact type). ∎
@@
In this section, we give a proof of Theorem 1.10.

### v4

@@ 2.4. The structure theorem of morphism
Let $(M,\omega_{g})$ be a compact Kähler manifold. Assume that $\mathcal{F}$ is an invertible sheaf of $\mathcal{O}(\Lambda^{p,0}M)$ which is pseudo-effective as a line bundle. Then there exist a finite covering $\{U_{\alpha}\}$ consists of coordinate neighborhoods and local holomorphic $(p,0)$-forms $\eta_{\alpha}$ generating $\mathcal{F}|_{U_{\alpha}}$, $\mathcal{F}$ is associated with the C̆ech cocycle $b_{\alpha\beta}$ in $\mathcal{O}_{M}^{*}(U_{\alpha})$ such that $\eta_{\beta}=b_{\alpha\beta}\eta_{\alpha}$ and there exists a singular Hermitian metric $h=e^{-\varphi}$ of $\mathcal{F}$ defined by a collection of plurisubharmonic functions $\varphi_{\alpha}\in\mathrm{PSH}(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$ such that $e^{-\varphi_{\beta}}=|b_{\alpha\beta}|^{2}e^{-\varphi_{\alpha}}$. We can define a globally bounded measurable function $\psi=e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}$ as in ([9]), which can be viewed as the hermitian metric ratio $\frac{|\eta_{\alpha}|^{2}_{g}}{|\eta_{\alpha}|^{2}_{h}}$. Since $\sqrt{-1}\partial\bar{\partial}$ of the plurisubharmonic function $\varphi_{\alpha}$ is a closed positive current, We have
$\sqrt{-1}\partial\bar{\partial}(e^{\varphi_{\alpha}})=e^{\varphi_{\alpha}}(\sqrt{-1}\partial\varphi_{\alpha}\wedge\bar{\partial}\varphi_{\alpha}+\sqrt{-1}\partial\bar{\partial}\varphi_{\alpha}),$
which is also a closed positive current on $U_{\alpha}$. This shows that $\partial\varphi_{\alpha}$ is $L^{2}$ with respect to the wight $e^{\varphi_{\alpha}}$ and $e^{\varphi_{\alpha}}\partial\bar{\partial}\varphi_{\alpha}$ has locally finite measure corfficients on $U_{\alpha}$. Hence
@@
This subsection is devoted to the proof of Proposition 3.1. Firstly, we will show the following key identity for forms.
@@
$\omega^{n-p-1}=(\sqrt{-1})^{n-p-1}(n-p-1)!\sum\limits_{i_{1}<i_{2}<\cdots<i_{n-p-1}}dz^{i_{1}}\wedge d\bar{z}^{i_{1}}\wedge\cdots\wedge dz^{i_{n-p-1}}\wedge d\bar{z}^{i_{n-p-1}}$
@@
$(\sqrt{-1})^{p^{2}}\alpha\wedge\frac{\eta\wedge\bar{\eta}}{p!}=$ $(\sqrt{-1})^{p+1}\sum\limits_{J_{p}}\sum\limits_{i\notin J_{p}}\alpha_{i\bar{i}}\eta_{J_{p}}\overline{\eta_{J_{p}}}dz^{i}\wedge d\bar{z}^{i}\wedge dz^{j_{1}}\wedge d\bar{z}^{j_{1}}\wedge\cdots\wedge dz^{j_{p}}\wedge d\bar{z}^{j_{p}}$
@@
where $\beta_{\alpha}=\Lambda^{p-1}((\sqrt{-1})^{p^{2}}\frac{\eta_{\alpha}\wedge\overline{\eta_{\alpha}}}{p!})$. $e^{\varphi_{\alpha}}\beta_{\alpha}$ and $e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}}$ are globally defined differential forms with bounded measurable cofficients. For clarity, the arguement will be divided into three steps.
@@
Although $\varphi_{\alpha}$ is in general singular, we can still perform integration by parts as in [13]. For the reader’s convenience, we provide a detailed explaination.
Consider a partitions of unity $\{\rho_{\alpha}\}$ subordinated to the finite cover $\{U_{\alpha}\}$. Since $\mathrm{supp}\rho_{\alpha}\subset\subset U_{\alpha}$, for any $\epsilon>0$ sufficiently small, there exists a sequence of smooth plurisubharmonic functions $\varphi_{\alpha,\epsilon}$ decreasing to $\varphi_{\alpha}$ on some open subset $V_{\alpha}$ such that $\mathrm{supp}\rho_{\alpha}\subset\subset V_{\alpha}\subset\subset U_{\alpha}$. Applying the results of Bedford-Taylor (Lemma 2.4) to the uniformly bounded function $e^{\varphi_{\alpha,\epsilon}}$ on $V_{\alpha}$, we obtain local weak covergence
@@
The last equality follows from the fact $\sum\limits_{\alpha}\rho_{\alpha}=1$. Combining (3.10) with (3.9), we obtain
@@
Next, we analyze each term on the right-hand side of integral inequality (3.11) seperately. First, we consider the sign of first term in (3.11).
@@
For any $x_{0}\in U_{\alpha}$, choose local holomorphic coordinates centered at $x_{0}$ such that
@@
at $x_{0}$. For simplicity, we omit the index $\alpha$ in the following calculation. We write $\eta=\sum\limits\eta_{I_{p}}dz^{I_{p}}$. Recall (3.3) and (3.7), which give us
@@
By direct calculation, we derive the following Bochner type formula in the sense of distribution,
@@
$\begin{split}e^{2\varphi}\langle\operatorname{Rm},-\beta\otimes\beta\rangle_{g}=e^{2\varphi}\sum\limits_{i,k}R_{i\bar{i}k\bar{k}}\beta_{i\bar{i}}\beta_{k\bar{k}}\geq&e^{2\varphi}\frac{\kappa(x_{0})}{2}\left[(\sum\limits_{i}\beta_{i\bar{i}})^{2}+\sum\limits_{i}\beta_{i\bar{i}}^{2}\right]\\ =&\frac{\kappa(x_{0})}{2}e^{2\varphi}\left(|\eta|_{g}^{4}+|\beta|_{g}^{2}\right)\\ \geq&\frac{\kappa(x_{0})}{2}\psi^{2}\end{split}$ (3.16)
@@
¯$\partial\bar{\partial}(e^{\varphi_{\alpha}}\eta_{\alpha}\wedge\overline{\eta_{\alpha}})=\partial(e^{\varphi_{\alpha}}\bar{\partial}\varphi_{\alpha}\wedge\eta_{\alpha}\wedge\overline{\eta_{\alpha}}+e^{\varphi_{\alpha}}(-1)^{p}\eta_{\alpha}\wedge\overline{\partial\eta_{\alpha}})$
@@
Hence $d\psi=0$ almost everywhere on $M$, which implies that $\psi=C$ for some constant $C\geq 0$ almost everywhere. We claim that $C>0$. Suppose not, we have $\psi=0$ almost everywhere on $U_{\alpha}$. Since $\varphi_{\alpha}\in Psh(U_{\alpha})\cap L_{loc}^{1}(U_{\alpha})$, $\{\varphi_{\alpha}=-\infty\}$ has zero Lebesgue measure. This fact implies that $\eta_{\alpha}\equiv 0$ on $U_{\alpha}$ due to the smoothness of $\eta_{\alpha}$, which contradicts the fact that $\eta_{\alpha}$ generates $\mathcal{F}|_{U_{\alpha}}$. Hence, $\varphi_{\alpha}<+\infty$ and $e^{\varphi_{\alpha}}|\eta_{\alpha}|^{2}_{g}=C>0$ implies that $|\eta_{\alpha}|^{2}_{g}>0$ and $\varphi_{\alpha}=\log C-\log|\eta_{\alpha}|^{2}_{g}$ is a smooth function on $U_{\alpha}$. The following calculation will be done in the smooth sense.
@@
Since $x_{0}\in M$ is arbitrary, we have ¯$\partial\bar{\partial}\varphi=0$ on $M$ and $D^{\prime}\eta+\eta\partial\varphi=0$ on $M$. The fact that ¯$\partial\bar{\partial}\varphi_{\alpha}=0$ implies that $(\mathcal{F},e^{-\varphi})$ is a flat line bundle. In each coordinate neighborhood $U_{\alpha}$, $\varphi_{\alpha}$ is pluriharmonic, so it can be written as $\varphi_{\alpha}=f_{\alpha}+\overline{f_{\alpha}}$ for some holomorphic function $f_{\alpha}\in\mathcal{O}(U_{\alpha})$. Therefore,
$D(e^{f_{\alpha}}\eta_{\alpha})=e^{f_{\alpha}}(D^{\prime}\eta_{\alpha}+\eta_{\alpha}\partial\varphi_{\alpha})=0$ (3.25)
where $e^{f_{\alpha}}\eta_{\alpha}$ is a local holomorphic section of $\mathcal{F}|_{U_{\alpha}}$. This shows that $\mathcal{F}$ is invariant by parallel transport induced by connection $D$ of $\Lambda^{p,0}M$, which verifies the statement (1).
We obtain a global real semi-positive $(1,1)$-form given by
$\beta:=e^{\varphi_{\alpha}}\beta_{\alpha}=\Lambda^{p-1}\left((\sqrt{-1})^{p^{2}}\frac{e^{f_{\alpha}\eta_{\alpha}}\wedge\overline{e^{f_{\alpha}}\eta_{\alpha}}}{p!}\right),$
which is determined by $\mathcal{F}$ and is parallel with respect to $D$. We then define a parallel and self-adjoint linear transformation $S:T^{1,0}M\rightarrow T^{1,0}M$ by
$\langle S(X),Y\rangle_{g}=\beta(X,\bar{Y}),\ \forall X,Y\in T^{1,0}M.$ (3.26)
From the above computation at a fixed point $x_{0}$, $V_{x_{0}}=\operatorname{span}\{\frac{\partial}{\partial z^{i}}\}_{i=1}^{m}$ and $W_{x_{0}}=\operatorname{span}\{\frac{\partial}{\partial z^{i}}\}$ consist of eigenvectors corresponding to nonzero and zero eigenvalues of $S_{x_{0}}$, respectively. We also have that every $v\in T_{x_{0}}^{1,0}M$ is truly flat and $\dim V_{x_{0}}=m>0$. Since $S$ is parallel, $x\mapsto V_{x}$ and $x\mapsto W_{x}$ define parallel distrubutions. The proof is complete. ∎
3.3. Proof of Theorem 1.5 and Theorem 1.7
In this subsection, we proceed to show Theorem 1.5 and Theorem 1.7. The proofs are based on Theorem 1.4, Campana-Demailly-Peternell’s critetion, de Rham’s splitting theorem ([17]).
@@
Since the holomorphic sectional curvature is nonnegative and there is no nonzero truly flat tangent vector at some point, by Theorem 1.4 any invertible subsheaf $\mathcal{F}\subset\mathcal{O}(\Lambda^{p,0}M)$ is not pseudo-effective. This actually implies that $H^{p,0}_{\bar{\partial}}(M)=0,\ \forall 1\leq p\leq n$. In particular, M is projective by Kodaira’s theorem ([25]). Combining this with Campana-Demailly-Peternell’s criterion (Lemma 2.5), we conclude that $M$ is rationally connected. ∎
@@
We will argue by contradiction. Suppose there exists an invertible subsheaf $\mathcal{F}$ of $\mathcal{O}(\Lambda^{p,0}M)$ that is pseudoeffective as a line bundle. According to Theorem 1.4, $T^{1,0}M$ admits a decomposition $V+W$ of $T^{1,0}M$ where $V$ and $W$ consist of eigenvectors corresponding to nonzero and zero eigenvalues of $S$, respectively, and $\forall p\in M$, every $v\in V_{p}$ is truly flat. Since $M$ is simply connected, by de Rham’s splitting theorem ([17, Theorem II]), $(M,\omega)\simeq(M_{1},\omega_{1})\times(M_{2},\omega_{2})$ isometrically and holomorphically, with $TM_{1}=V$. This implies that $M_{1}$ is a simply connected Kähler manifold with the flat metric $\omega_{1}$. According to the classification of complex space forms, $M_{1}$ must be $\mathbb{C}^{m}$ with $m=\dim V>0$, which contradicts the fact that $M$ is compact. By Campana-Demailly-Peternell’s criterion (Lemma 2.5), the proof is complete. ∎
@@
In this section, we provide a proof of Theorem 1.9. The proof relies on $L^{2}$-vanishing theorem (Proposition 4.1), Atiyah’s $L^{2}$-index Theorem ([1]) and properties of Albanese maps. First, we present the following $L^{2}$-vanishing theorem for nonnegative holomorphic sectional curvature.
Proposition 4.1.
Let $(M,\omega)$ be a complete Kähler manifold with nonnegative holomorphic sectional curvature, then any $L^{2}$-holomorphic $(p,0)$-form is parallel.
The idea of proof is essentially the same as that for Theorem 1.4, but with some minor differences, because $|\eta|^{4}_{g}$ may be not $L^{1}$-integrable. For the reader’s convenience, we provide a brief proof in the appendix. From Proposition 4.1 and Atiyah’s $L^{2}$-index theorem ([1]), we derive the following proposition.
Proposition 4.2.
Let $(M,\omega)$ be a compact Kähler manifolds with nonnegative holomorphic sectional curvature. One of the two following situations occurs:
  • (i)
    $\pi_{1}(M)$ is finite, then $M$ is projective and rationally connected.
  • (ii)
    $\pi_{1}(M)$ is infinite, then $\chi(M,\mathcal{O}_{X})=0$.
Proof.
Let $f:\widetilde{M}\rightarrow M$ be the universal cover of $M$, then $(\widetilde{M},f^{*}\omega)$ is a compact simple connected Kähler manifold with nonnegative holomorphic sectional curvature. If $\pi_{1}(M)$ is finite, then $(\widetilde{M},f^{*}\omega)$ is compact simple connected, hence projective and rationally connected (Theorem 1.7). Since a pseudoeffective invertible subsheaf $\mathcal{F}$ of $\mathcal{O}(\Lambda^{p,0}M)$ induces a pseudoeffective invertible subsheaf $f^{*}\mathcal{F}$ of $\mathcal{O}(\Lambda^{p,0}\widetilde{M})$, $M$ must be projective and rationally connected by the Compana-Demailly-Peternell’s criterion (Lemma 2.5).
If $\pi_{1}(M)$ is infinite, then $(\widetilde{M},f^{*}\omega)$ is noncompact and complete. Denote the Hilbert space of $L^{2}$-harmonic form of degree $(p,q)$ by $H^{p,q}_{L^{2}}(\widetilde{M})$. According to $L^{2}$-Hodge theory (see [20]), a $L^{2}$ $(p,0)$-form is harmonic if and only if it is a holomorphic $(p,0)$-form. In general, $H^{p,q}_{L^{2}}(\widetilde{M})$ might be infinite-dimensional. Using the isometric action of $\Gamma:=\pi_{1}(X)$ on them, one can associate to them a nonnegative real number $\dim_{\Gamma}(H^{p,q}_{L^{2}}(\widetilde{M}))$ (c.f. [1]). $\dim_{\Gamma}(H^{p,q}_{L^{2}}(\widetilde{M}))=0$ if and only if $H^{p,q}_{L^{2}}(\widetilde{M})=\{0\}$. By Atiyah’s $L^{2}$-index theorem (c.f. [1, 20]), we know that
$\chi(M,\mathcal{O}_{M})=\chi_{L^{2}}(\widetilde{M},\mathcal{O}_{\widetilde{M}}):=\sum\limits_{k=0}^{\dim M}(-1)^{k}\dim_{\Gamma}(H^{0,q}_{L^{2}}(\widetilde{M})).$
By Proposition 4.1, the norm $|\eta|_{f^{*}g}$ of any $L^{2}$-holomorphic $(p,0)$-form $\eta$ is equal to some constant. Because the volume of $(\widetilde{M},f^{*}\omega)$ is infinite, any $L^{2}$-holomorphic $(p,0)$-form must be zero. Thus $H^{p,0}_{L^{2}}(M)=\{0\}$ for every $0\leq p\leq n$. By conjugation, $H^{0,p}_{L^{2}}(M)=\{0\}$ for every $0\leq p\leq n$. Consequently, $\chi(M,\mathcal{O}_{M})=0$. ∎
From Proposition 4.1, we have the following fact about the Albanese map.
Proposition 4.3.
Let $(M,\omega)$ be a compact Kähler manifold with nonnegative holomorphic sectional curvature. Then the Albanese map $\alpha:=M\rightarrow A(M)$ is a submersion. In particular, it is surjective.
Proof.
Let $q=\dim H^{1,0}_{\bar{\partial}}(M)$ and let $\{\eta_{1},\cdots,\eta_{q}\}$ be a basis of holomorphic $(1,0)$-forms. Recall that the Albanese map $\alpha$ is defined by
$x\mapsto\left(\int_{a}^{x}\eta_{1},\cdots,\int_{a}^{x}\eta_{q}\right)\mod\Lambda,$
where $\Lambda\subset\mathbb{C}^{q}$ (see, e.g., [14, Chapter VI, section 9.2]). Then $d\alpha=(\eta_{1},\cdots,\eta_{q})$ at every point. If $\operatorname{rank}(d\alpha)<q=\dim(A(M))$ at some point $x$, then there would exist a non-zero linear combination
$u=\lambda_{1}\eta_{1}+\cdots+\lambda_{q}\eta_{q}$
with $u(x)=0$. By Proposition 4.1, $u$ must be zero, which leads to a contradiction. ∎
Now let us proceed to show Theorem 1.9.
Proof of Theorem 1.9.
Since $M$ is non-projective, it follows from Kodaira’s theorem ([25]) that $M$ admits a nonzero holomorphic $(2,0)$-form, and from Proposition 4.2 that $\chi(M,\mathcal{O}_{M})=0$. Let $h^{p,0}$ denote $\dim H^{p,0}_{\bar{\partial}}(M)$.
If $h^{3,0}>0$, there admits a nonzero holomorphic $(3,0)$-form $\eta$. By Proposition 4.1, $\eta$ must be parallel and therefore doesn’t vanish anywhere. Since $\dim M=3$, the existence of the nowhere-vanishing holomorphic $(3,0)$-form $\eta$ implies that $K_{M}\simeq\mathcal{O}_{M}$. Consequently, $K_{M}$ is flat, and we have:
$0=-n\int_{M}c_{1}(K_{M})\wedge\omega^{n-1}=n\int_{M}\operatorname{Ric}(\omega)\wedge\omega^{n-1}=\int_{M}S\cdot\omega^{n},$
where $S$ is the scalar curvature of $\omega$. According to Berger’s averaging trick ([3]),
[…]
at any point $x$ of $M$. Hence, the holomorphic sectional curvature is identically zero, and thus $M$ is flat.
If $h^{3,0}=0$, then $01-h^{1,0}+h^{2,0}=\chi(M,\mathcal{O}_{M})=$ 0, which implies that $h^{1,0}=1+h^{2,0}\geq 2$. Suppose that $h^{1,0}=3$. Then we have $h^{3,0}\geq 1$, which is a contradiction. Therefore, $h^{1,0}=2$. From Matsumura’s result (Lemma 2.9) and Proposition 4.3, the Albanese map $\alpha:M\rightarrow\mathbb{T}^{2}$ is locally trivial and smooth. The fibre $F$ of $\alpha$ admits a Kähler metric $\omega_{F}$ with nonnegative holomorphic sectional curvature, and the universal cover $(\widetilde{M},\widetilde{\omega})\cong(\widetilde{F},\widetilde{\omega_{F}})\times(\mathbb{C}^{2},\omega_{0}),$ where $(\widetilde{M},\widetilde{\omega})$ and $(\widetilde{F},\widetilde{\omega_{F}})$ are the universal covers of $(M,\omega)$ and $(F,\omega_{F})$ respectively, $\omega_{0}$ is the flat metric. Since $\dim F=1$, the holomorphic sectional curvature of $\omega_{F}$ is either identically zero or quasi-positive. Hence, either $M$ is flat, or $F$ is $\mathbb{P}^{1}$. When $F$ is $\mathbb{P}^{1}$, $M$ is a $\mathbb{P}^{1}$-bundle over a torus $\mathbb{T}^{2}$, and the universal cover is $\mathbb{P}^{1}\times\mathbb{C}^{2}$. ∎
5. Proof of Theorem 1.12
In this section, we provide a proof of Theorem 1.12.